OpenStudy (anonymous):

Whats the correct simplification of (8x + 5)(4x - 6).

OpenStudy (anonymous):

32x^2 -28x-30 :D

OpenStudy (callisto):

(8x + 5)(4x - 6) = 8x (4x - 6) +5 (4x - 6) = 32x^2 - 48x + 20x -30 =?

OpenStudy (anonymous):

hope it helps!:)

OpenStudy (callisto):

@seashell Only giving answer won't help much. Also, you've violated the Code of Conduct..

OpenStudy (anonymous):

D: sorry..